OpenStudy (anonymous):

How do you solve this? 14² + 1/2 · 14 · 8 + 1/2(3.14)(7²).

OpenStudy (anonymous):

is there x's in the equation or just those numbers

OpenStudy (anonymous):

It has a picture of a triangle and a square (triangle has a side height of 8" and the square with a height of 14)

OpenStudy (anonymous):

The question asks "Which is the area closest to the area of the wall hanging?"

OpenStudy (anonymous):

14² = 196 1/2*14*8 = 56 1/2(3.14)(7²) = 76.93 196 + 56 + 76.93 = 328.93 thats the answer if i understand the question correctly

OpenStudy (anonymous):

Thank you it's one of the answer choices , I get it now

OpenStudy (anonymous):

what class is that?

OpenStudy (anonymous):

Algebra 1